A pharmaceutical company receives large shipments of aspirin tablets. The acceptance sampling plan is to randomly select and test 41 ​tablets, then accept the whole batch if there is only one or none that​ doesn't meet the required specifications. If one shipment of 7000 aspirin tablets actually has a 33​% rate of​ defects, what is the probability that this whole shipment will be​ accepted? Will almost all such shipments be​ accepted, or will many be​ rejected?

The probability that this whole shipment will be accepted is___?

The company will accept ___% and reject ___%
Answer #1

Answer:

Given,

Sample n = 41

p = 3% = 0.03

q = 1 - 0.03

= 0.97

Probability of zero defects = nC0*p^r*q^(n-r)

= 41C0*0.03^0*0.97^41

= 0.97^41

= 0.2868

P(1) = nC1*p^r*q^(n-r)

= 41C1*0.03*(0.97^40)

= 0.3637

Required probability = P(0) + P(1)

= 0.2868 + 0.3637

= 0.6505

= 65.05% is accepted & remaining 34.95% will rejected.